1. A designer, attempting to arrange the characters of his artwork in the form of a square grid with equal number of rows and columns, found that 24 characters were left out. When he tried to add one more row and column, he found that he was short of 25 characters. Can you find the number of characters used by the designer?: 1. A designer, attempting to arrange the characters of his artwork in the form of a square grid with equal number of rows and columns, found that 24 characters were left out. When he tried to add one more row and column, he found that he was short of 25 characters. Can you find the number of characters used by the designer?
Answer by ankor@dixie-net.com(4182)

"A designer, attempting to arrange the characters of his artwork in the form of a square grid with equal number of row and columns, found that 24 characters were left out."
:
If x = number of characters in each row
we translate this statement as
x*x + 24
x^2 + 24 = total number of characters
:
"When he tried to add one more row and column, he found that he was short of 25 characters."
:
Now, the number of rows is "x + 1", and the number of columns is "x + 1", so we have
(x + 1)(x + 1) - 25
FOIL
x^2 + 2x + 1 - 25 = total number of charaters
:
an equation using both statement:
x^2 + 2x + 1 -25 = x^2 + 24
x^2 + 2x -24 = x^2 + 24
x^2 - x^2 + 2x - 24 - 24 = 0
2x - 48 = 0
2x = +48
x = 48/2
x = 24
:
Grid is 24 by 24 which contains 576 character + 24 left out = 600
Grid of 25 by 25 = 625 character so he would have 25 blanks